# Theoretical Insights into Fine-Tuning Attention Mechanism: Generalization and Optimization

### v1

@@ 6 Conclusion and Limitation
In this paper, we present our key findings in fine-tuning attention mechanism: Different Impact—optimizing the $\mathbf{W}_{v}$ matrix significantly improves performance compared to the $\mathbf{W}_{k}$ matrix. Fine-tuning only the $\mathbf{W}_{q}$ and $\mathbf{W}_{v}$ matrices is computationally efficient and can yield results that match or surpass fine-tuning all three matrices $\mathbf{W}_{q}$, $\mathbf{W}_{k}$, and $\mathbf{W}_{v}$. Efficient Convergence—using distinct learning rates for these matrices is essential for optimal performance, with a higher learning rate for $\mathbf{W}_{v}$ speeding up convergence. While theoretical analysis of these phenomena is limited, this paper provides insights from two angles: Generalization—fine-tuning only $\mathbf{W}_{q}$ and $\mathbf{W}_{v}$ improves generalization and memory efficiency, and Optimization—using different learning rates enhances the efficiency of feature learning in the attention mechanism, leading to more effective fine-tuning. Our analysis provides a theoretical foundation for the configuration and improvement of lightweight algorithms in LLMs fine-tuning. However, further studies are required on (i) how the task and neural architecture influence the optimal ratio $\lambda$ and (ii) whether the results about attention hold true for tasks beyond natural language processing. These studies will further deepen our understanding of attention-based fine-tuning in LLMs.
Appendix A More Related Works
Attention mechanism analysis. A key component of transformers is the attention mechanism, which dates back to [7]. Initially designed to capture long-range signals in sequential inputs by mixing individual tokens, it has also been utilized to capture general structures in input data. After the fully-attention-based language model has appeared [39, 4], the research community gets interested in the functionality and benefits of the attention. For instance, transformers implicitly favor hierarchical interpretations of input sequences [19], the computational graphs tend to be tree-structured [27, 29]. Theoretical analysis of training dynamics sheds light on how to identify key tokens [37], select a few relevant tokens only (which is called localized attention) or select many tokens uniformly [3], and learn topic structure [23]. Besides, considerable works [32, 1, 47] try to understand in-context learning capabilities from the perspective of gradient descent with attention.
Scaling for neural networks. Scaling refers to the process of enlarging a specific ingredient of a model to enhance its overall performance [16]. The method is straightforward: extend the width or depth of a neural network towards infinity, analyze how this limit is influenced by hyperparameters like the learning rate and initialization variance during training, and then establish well-founded choices for these hyperparameters to achieve a specific objective [15, 34, 9, 43, 45, 12, 8, 46, 11]. In the theory of scaling of neural networks, one usually tracks the asymptotic behaviour of key quantities as we scale some model ingredient, it is a standard approach used to derive scaling rules for initialization [34], activation function [9], network parametrization [46]. In this paper, we are interested in scaling model capacity via the width n for the fact that most state-of-the-art pre-trained models have large width. Examples of the infinite-width limit can be found in studies focused on initialization methods [15, 43], or more comprehensive approaches to network parameterization. For instance, Yang et al. [45] introduced µP, a parameterization technique for neural networks that guarantees feature learning in the infinite-width limit, providing specific scaling rules for both architecture and learning rates to optimize feature learning [44, 45].
Parameter-efficient fine-tuning. Fine-tuning all the parameters of a large language models, known as full fine-tuning, is highly computationally expensive. To reduce the computational cost, various parameter-efficient fine-tuning (PEFT) methods have been proposed [5], which only fine-tune a small number of (extra) model parameters. PEFT methods can be divided into two categories from the perspective of whether extra parameters are involved: (1) extra-parameter methods, freeze all of the original parameters of an LLM and insert a set of learnable parameters to optimize the model input or model layers suach as adapter tuning [17] , prompt tuning [21] and prefix tuning [22]; (2) intra-parameter methods freeze most of the original parameters of an LLM and only tune a small number of parameters of the LLM such as LoRA [18]. Furthermore, He et al. [14] present a unified framework that establishes connections between PEFT methods and Zhu et al. [48] formally identify and investigate asymmetry in the roles of low-rank adapter matrices in LoRA fine-tuning.
Appendix B Omitted Proofs and Additional Results
B.1 The connection between Prefix tuning and LoRA.
@@ Appendix B Omitted Proofs and Additional Results
$=(1-\lambda(\mathbf{x}))\text{softmax}(\mathbf{x}\mathbf{W}_{q}\mathbf{W}_{k}^{T}\mathbf{C}^{T})\mathbf{C}\mathbf{W}_{v}+\lambda(\mathbf{x})\text{softmax}(\mathbf{x}\mathbf{W}_{q}\mathbf{P}_{k}^{T})\mathbf{P}_{v}$
$=(1-\lambda(\mathbf{x}))\overbrace{\text{Attn}(\mathbf{x}\mathbf{W}_{q},\mathbf{CW}_{k},\mathbf{CW}_{v})}^{{\color[rgb]{0,0,1}\text{standard attention}}}+\lambda(\mathbf{x})\overbrace{\text{Attn}(\mathbf{x}\mathbf{W}_{q},\mathbf{P}_{k},\mathbf{P}_{v})}^{{\color[rgb]{0.0,0.5,0.0}\text{independent of C}}},$
where $\lambda(\mathbf{x})=\frac{\sum_{i}\text{exp}(\mathbf{xW}_{q}\mathbf{P}_{k}^{T})_{i}}{\sum_{i}\text{exp}(\mathbf{xW}_{q}\mathbf{P}_{k}^{T})_{i}+\sum_{j}\text{exp}(\mathbf{xW}_{q}\mathbf{W}_{k}^{T}\mathbf{C}^{T})_{j}}$ is a scalar that represents the sum of normalized attention weights on the prefixes. Notice that the first term in blue represents the original attention mechanism without prefixes, while the second term in green introduces a position-wise adjustment independent of $\mathbf{C}$. It provides an alternative perspective on Prefix tuning, where a position-wise modification is applied to the original attention output $\mathbf{h}$ via linear interpolation:
$\mathbf{h}\leftarrow(1-\lambda(\mathbf{x})\mathbf{h}+\lambda(\mathbf{x})\Delta\mathbf{h},\ \ \ \Delta\mathbf{h}:=\text{softmax}(\mathbf{xW}_{q}\mathbf{P}_{k}^{T})\mathbf{P}_{v}\triangleq\text{softmax}(\mathbf{xA})\mathbf{B}.$
B.2 Proof of Theorem 1
@@
(Xu and Raginsky [42, Theorem 1.]). Assume the loss $\ell(\mathbf{W},Z)$ is R-subGaussian for any $\mathbf{W}\in\mathcal{W}$, then
@@
B.3 Initialization Discussion
Following standard initialization schemes (e.g., LeCun Init and He Init [20, 15]), one generally consider a Gaussian initialization of the weights as follows: $a_{i}\sim\mathcal{N}(0,\sigma_{a}^{2}),\quad b\sim\mathcal{N}(0,\sigma_{b}^{2})$ (The Gaussian distribution can be substituted with any other distribution that has finite variance). Revisiting Starting with a Toy setting, $a\in\mathbb{R}^{1\times n},b\in\mathbb{R}$. Thus, one should set $\sigma_{a}^{2}=\Theta(n^{-1}),\sigma_{b}^{2}=0$ to ensure $xa^{T}$ does not explode with width ($xa^{T}=\Theta(1)$), for a non-zero initialization for $a$. This is justified by the Central Limit Theorem (See [45] for more technical details). And if we choose a non-zero initialization for $b$, one should make sure that $\sigma_{b}^{2}=\Theta(1),\sigma_{a}^{2}=0$. And we will consider these two initialization schemes to show our theoretical understanding.
B.4 Gamma Function
@@
In Section 4.2, the learning rate $\eta_{a}=\Theta(n^{c_{a}}),\eta_{b}=\Theta(n^{c_{b}})$ for some $c_{a},c_{b}\in\mathbb{R}$. And in Appendix B.3 we assume that the init weights are also scale polynomially with n, it is evident that preactivations, gradients, and weight updates all exhibit asymptotic polynomial growth in n.
@@
  • Addition. Generally, we have $\gamma[v_{1}+v_{2}]=\max(\gamma[v_{1}],\gamma[v_{2}])$. The only instance where this does not hold is when $v_{1}=-v_{2}$. This is typically a zero-probability event if the two variables are random variables that are not perfectly correlated, which is the case in most scenarios where we apply this formula (Appendix B.5).
B.5 Proof of Theorem 2
@@
In Section 3, we say that the feature learning of attention mechanism is efficient when $\delta_{t}^{i}=\Theta(1)$ for all $t,i\in\{1,2,3\}$. Using the elementary formulas from Appendix B.4, we can get (for all $t$):
@@
Starting from $t=1$. In both initialization schemes discussed in Appendix B.3, we have to set $\gamma[\eta_{b}]=0$ and $\gamma[\eta_{a}]=-1$ to ensure that $\gamma[f_{t}]=\gamma[xa_{t}^{T}]+\gamma[b_{t}]=0$:
@@
Appendix C Extension to Experiments
C.1 Empirical Details
C.1.1 GLUE Tasks with RoBERTa
@@
C.1.2 Training Hyperparameters
Training hyperparameters.
@@ C.1.2 Training Hyperparameters
C.1.3 MNLI Task with Llama3.1-8b
@@
C.2 Empirical Results
C.2.1 GLUE Tasks Train Loss
@@
C.2.3 Algorithm Framework

### v3

@@ 6 Conclusion and Limitation
In this paper, we present our key findings in fine-tuning attention mechanism: Unequal Importance of Attention Matrices and Attention Matrices with Customized Learning Rate Lead to Better Convergence. While theoretical analysis of these phenomena is limited, this paper provides insights from two angles: Generalization—fine-tuning only $\mathbf{W}_{q}$ and $\mathbf{W}_{v}$ improves generalization and memory efficiency, and Optimization—using different learning rates enhances the efficiency of feature learning in the attention mechanism, leading to more effective fine-tuning. Our analysis provides a theoretical foundation for the configuration and improvement of lightweight algorithms in LLMs fine-tuning. However, further studies are required on (i) how task type and architecture affect the optimal learning rate ratio $\lambda$, and (ii) whether these findings extend beyond NLP tasks. These studies will further deepen our understanding of attention-based fine-tuning.
Ethical Statement
There are no ethical issues.
Acknowledgments
We sincerely appreciate the anonymous reviewers for their valuable suggestions. This research was supported by National Natural Science Foundation of China (No.62476277), National Key Research and Development Program of China (NO. 2024YFE020320), CCF-ALIMAMA TECH Kangaroo Fund(No.CCF-ALIMAMA OF 2024008), and Huawei-Renmin University joint program on Information Retrieval. We also acknowledge the support provided by the fund for building worldclass universities (disciplines) of Renmin University of China and by the funds from Beijing Key Laboratory of Big Data Management and Analysis Methods, Gaoling School of Artificial Intelligence, Renmin University of China, from Engineering Research Center of Next-Generation Intelligent Search and Recommendation, Ministry of Education, from Intelligent Social Governance Interdisciplinary Platform, Major Innovation & Planning Interdisciplinary Platform for the “DoubleFirst Class” Initiative, Renmin University of China, from Public Policy and Decision-making Research Lab of Renmin University of China, and from Public Computing Cloud, Renmin University of China.
Appendix A Omitted Proofs and Additional Results
A.1 The connection between Prefix tuning and LoRA.
@@ Appendix A Omitted Proofs and Additional Results
$=(1-\alpha(\mathbf{x}))\text{softmax}(\mathbf{x}\mathbf{W}_{q}\mathbf{W}_{k}^{T}\mathbf{C}^{T})\mathbf{C}\mathbf{W}_{v}+\alpha(\mathbf{x})\text{softmax}(\mathbf{x}\mathbf{W}_{q}\mathbf{P}_{k}^{T})\mathbf{P}_{v}$
$=(1-\alpha(\mathbf{x}))\overbrace{\text{Attn}(\mathbf{x}\mathbf{W}_{q},\mathbf{CW}_{k},\mathbf{CW}_{v})}^{{\color[rgb]{0,0,1}\text{standard attention}}}+\alpha(\mathbf{x})\overbrace{\text{Attn}(\mathbf{x}\mathbf{W}_{q},\mathbf{P}_{k},\mathbf{P}_{v})}^{{\color[rgb]{0.0,0.5,0.0}\text{independent of C}}},$
where $\alpha(\mathbf{x})=\frac{\sum_{i}\text{exp}(\mathbf{xW}_{q}\mathbf{P}_{k}^{T})_{i}}{\sum_{i}\text{exp}(\mathbf{xW}_{q}\mathbf{P}_{k}^{T})_{i}+\sum_{j}\text{exp}(\mathbf{xW}_{q}\mathbf{W}_{k}^{T}\mathbf{C}^{T})_{j}}$ is a scalar that represents the sum of normalized attention weights on the prefixes. Notice that the first term in blue represents the original attention mechanism without prefixes, while the second term in green introduces a position-wise adjustment independent of $\mathbf{C}$. It provides an alternative perspective on Prefix tuning, where a position-wise modification is applied to the original attention output $\mathbf{h}$ via linear interpolation:
$\mathbf{h}\leftarrow(1-\alpha(\mathbf{x})\mathbf{h}+\alpha(\mathbf{x})\Delta\mathbf{h},\ \ \ \Delta\mathbf{h}:=\text{softmax}(\mathbf{xW}_{q}\mathbf{P}_{k}^{T})\mathbf{P}_{v}\triangleq\text{softmax}(\mathbf{xA})\mathbf{B}.$
A.2 Proof of Theorem 1
@@
(Xu and Raginsky [2017, Theorem 1.]). Assume the loss $\ell(\mathbf{W},Z)$ is R-subGaussian for any $\mathbf{W}\in\mathcal{W}$, then
@@
A.3 Initialization Discussion
Following standard initialization schemes (e.g., LeCun Init and He Init [LeCun et al., 2002; He et al., 2016]), one generally consider a Gaussian initialization of the weights as follows: $a_{i}\sim\mathcal{N}(0,\sigma_{a}^{2}),\quad b\sim\mathcal{N}(0,\sigma_{b}^{2})$ (The Gaussian distribution can be substituted with any other distribution that has finite variance). Revisiting Starting with a Toy setting, $a\in\mathbb{R}^{1\times n},b\in\mathbb{R}$. Thus, one should set $\sigma_{a}^{2}=\Theta(n^{-1}),\sigma_{b}^{2}=0$ to ensure $xa^{T}$ does not explode with width ($xa^{T}=\Theta(1)$), for a non-zero initialization for $a$. This is justified by the Central Limit Theorem (See [Yang et al., 2022] for more technical details). And if we choose a non-zero initialization for $b$, one should make sure that $\sigma_{b}^{2}=\Theta(1),\sigma_{a}^{2}=0$. And we will consider these two initialization schemes to show our theoretical understanding.
A.4 Gamma Function
@@
In Section 4.2, the learning rate $\eta_{a}=\Theta(n^{c_{a}}),\eta_{b}=\Theta(n^{c_{b}})$ for some $c_{a},c_{b}\in\mathbb{R}$. And in Appendix A.3 we assume that the init weights are also scale polynomially with n, it is evident that preactivations, gradients, and weight updates all exhibit asymptotic polynomial growth in n.
@@
  • Addition. Generally, we have $\gamma[v_{1}+v_{2}]=\max(\gamma[v_{1}],\gamma[v_{2}])$. The only instance where this does not hold is when $v_{1}=-v_{2}$. This is typically a zero-probability event if the two variables are random variables that are not perfectly correlated, which is the case in most scenarios where we apply this formula (Appendix A.5).
A.5 Proof of Theorem 2
@@
In Section 3, we say that the feature learning of attention mechanism is efficient when $\delta_{t}^{i}=\Theta(1)$ for all $t,i\in\{1,2,3\}$. Using the elementary formulas from Appendix A.4, we can get (for all $t$):
@@
Starting from $t=1$. In both initialization schemes discussed in Appendix A.3, we have to set $\gamma[\eta_{b}]=0$ and $\gamma[\eta_{a}]=-1$ to ensure that $\gamma[f_{t}]=\gamma[xa_{t}^{T}]+\gamma[b_{t}]=0$:
@@
Appendix B Extension to Experiments
B.1 Empirical Details
B.1.1 GLUE Tasks with RoBERTa
@@
Training Hyperparameters. Training hyperparameters.
@@
B.1.2 MNLI Task with Llama3.1-8b
@@
Before fine-tuning model performance. QQP: 55.08 , MNLI: 33.34
B.2 More Discussions
Why fine-tune $\mathbf{W}_{q}\&\mathbf{W}_{v}$ instead of $\mathbf{W}_{k}\&\mathbf{W}_{v}$. In Eq.(1), the conventional attention function includes a term $\mathbf{xW}_{q}\mathbf{W}_{k}^{T}$. (1) In linear algebra, two matrices multiplied without an intermediate activation can be equivalent to a single matrix. Therefore, the effects of fine-tuning $\mathbf{W}_{q}\&\mathbf{W}_{v}$ and $\mathbf{W}_{k}\&\mathbf{W}_{v}$ are theoretically expected to yield similar outcomes (See the supplementary experimental results provided in Appendix x). (2) $\mathbf{W}_{k}$ operates only on the transformed representation matrix $\mathbf{xW}_{q}$ produced by the preceding transformation. Consequently, it loses direct access to the original representation information. This observation is consistent with the findings in [Qin et al., 2024]: the information representation in LoRA also exhibits significant limitations, as in $\Delta\mathbf{h}=\mathbf{xAB}$, where $\mathbf{B}$ similarly lacks access to the original representation information.
How to set the ratio $\lambda$? Naturally, as discussed in Remark 4, the optimal ratio $\lambda$ depends on the architecture and the fine-tuning task via the constants in $\Theta$ in Theorem 2. This is a limitation of these asymptotic results since they do not offer any insights on how the constants are affected by the task and the neural architecture. However, we can still employ some heuristic methods, such as: we can select an appropriate range by conducting a certain amount of experiments, as shown in Figure 1, it seems that a ratio of order $2^{1}-2^{4}$ is optimal. Moreover, $\lambda$ should not be too large; otherwise, as shown in the MNLI subplot in Figure 1, the model’s performance will collapse.
B.3 Empirical Results
GLUE Tasks Train Loss.
@@ B.3 Empirical Results
B.3.2 Ablation Experiments on Mistral-7B
In alignment with the experimental setup (hyperparameter setting for Llama3.1-8b) described in our Section 4.1, Figure 2, we have evaluated the RTE and MNLI task performances of our approach on Mistral-7B:
Method Hyperparameter RTE MNLI
LoRA (QKV) $r=16$, $\lambda=1$ 81.28 87.80
LoRA (QV) $r=16$, $\lambda=1$ 80.51 88.87
LoRA (QV) $r=16$, $\lambda=2$ 81.59 89.04
LoRA (QV) $r=16$, $\lambda=4$ 80.87 88.64
LoRA (QV) $r=16$, $\lambda=8$ 83.75 88.78
B.3.3 Algorithm Framework
@@
B.3.4 More Challenging Evaluation
Evaluating the model on more challenging benchmarks is essential for a comprehensive understanding of its capabilities. To address this, we follow [Yu et al., 2024] to fine-tune the LLaMA3.1-8B model on the MetaMathQA [Yu et al., 2024] dataset (the training set consists of the first 10K samples selected from the 150K MetaMathQA dataset.) and evaluate the performance on the GSM8K [Cobbe et al., 2021] (a benchmark for mathematical problem-solving).
Method GSM8K (100%)
Before fine-tune 25.55
LoRA (QKV) $r=16,\lambda=1$ 57.70
LoRA (QV) $r=16,\lambda=2$ 59.15
LoRA (QV) $r=16,\lambda=4$ 58.23
B.3.5 Fine-tuning K,V
We fine-tune only $\mathbf{W}_{k}$ and $\mathbf{W}_{v}$ of Roberta-base, others are the same to Table 1.
Weight Type $\mathbf{W}_{k},\mathbf{W}_{v}$ $\mathbf{W}_{q},\mathbf{W}_{v}$ $\mathbf{W}_{q},\mathbf{W}_{k},\mathbf{W}_{v}$
SST2(R) $r=8$ $0.920$ $0.919$ $0.922$
$r=16$ $0.920$ $0.921$ $0.923$
QNLI(R) $r=8$ $0.887$ $0.889$ $0.895$
$r=16$ $0.888$ $0.890$ $0.890$
QQP(R) $r=8$ $0.840$ $0.840$ $0.844$
$r=16$ $0.840$ $0.839$ $0.844$
MNLI(R) $r=8$ $0.821$ $0.820$ $0.827$
$r=16$ $0.822$ $0.824$ $0.828$
B.3.6 Directly fine-tuning Q,K,V with lambda
We fine-tuning $\mathbf{W}_{q},\mathbf{W}_{k},\mathbf{W}_{v}$ with $\lambda$ directly with the same settings in Table 2 for easy comparison, supporting one of our major claims in Theorem 2.
Method Trainable #Param (M) RTE STS-B MRPC CoLA
Before Fine-tune 0 45.12 -3.18 66.66 1.09
LoRA (QKV) $r=8,\lambda=1$ 1.62 70.76 90.25 85.04 58.03
LoRA (QKV) $r=8,\lambda=2$ 1.62 72.92 90.54 86.76 58.28
LoRA (QKV) $r=8,\lambda=4$ 1.62 73.64 90.84 87.74 60.66
LoRA (QKV) $r=8,\lambda=8$ 1.62 76.10 91.00 88.48 60.59
LoRA (QKV) $r=16,\lambda=1$ 2.07 70.39 90.25 86.03 58.04
LoRA (QKV) $r=16,\lambda=2$ 2.07 72.56 90.36 86.27 59.81
LoRA (QKV) $r=16,\lambda=4$ 2.07 74.00 90.84 86.76 60.07
LoRA (QKV) $r=16,\lambda=8$ 2.07 76.97 90.81 87.74 60.34
Appendix C More Related Works
Attention mechanism analysis. A key component of transformers is the attention mechanism, which dates back to [Graves, 2014]. Initially designed to capture long-range signals in sequential inputs by mixing individual tokens, it has also been utilized to capture general structures in input data. After the fully-attention-based language model has appeared [Vaswani et al., 2017; Brown et al., 2020], the research community gets interested in the functionality and benefits of the attention. For instance, transformers implicitly favor hierarchical interpretations of input sequences [Kharitonov and Chaabouni, 2021], the computational graphs tend to be tree-structured [Meng et al., 2022; Murty et al., 2023]. Theoretical analysis of training dynamics sheds light on how to identify key tokens [Tian et al., 2023], select a few relevant tokens only (which is called localized attention) or select many tokens uniformly [Bao et al., 2024], and learn topic structure [Li et al., 2023]. Besides, considerable works [Ren and Liu, 2023; Ahn et al., 2023; Yao et al., 2024] try to understand in-context learning capabilities from the perspective of gradient descent with attention.
Scaling for neural networks. Scaling refers to the process of enlarging a specific ingredient of a model to enhance its overall performance [Hoffmann et al., 2022]. The method is straightforward: extend the width or depth of a neural network towards infinity, analyze how this limit is influenced by hyperparameters like the learning rate and initialization variance during training, and then establish well-founded choices for these hyperparameters to achieve a specific objective [He et al., 2016; Schoenholz et al., 2017; Hayou et al., 2019; Yang, 2020; Yang et al., 2022; He et al., 2023; Hayou, 2023; Yang et al., 2023; Hayou et al., 2024b]. In the theory of scaling of neural networks, one usually tracks the asymptotic behaviour of key quantities as we scale some model ingredient, it is a standard approach used to derive scaling rules for initialization [Schoenholz et al., 2017], activation function [Hayou et al., 2019], network parametrization [Yang et al., 2023]. In this paper, we are interested in scaling model capacity via the width n for the fact that most state-of-the-art pre-trained models have large width. Examples of the infinite-width limit can be found in studies focused on initialization methods [He et al., 2016; Yang, 2020], or more comprehensive approaches to network parameterization. For instance, Yang et al. [2022] introduced µP, a parameterization technique for neural networks that guarantees feature learning in the infinite-width limit, providing specific scaling rules for both architecture and learning rates to optimize feature learning [Yang and Hu, 2021; Yang et al., 2022].
Parameter-efficient fine-tuning. Fine-tuning all the parameters of a large language models, known as full fine-tuning, is highly computationally expensive. To reduce the computational cost, various parameter-efficient fine-tuning (PEFT) methods have been proposed [Ding et al., 2023], which only fine-tune a small number of (extra) model parameters. PEFT methods can be divided into two categories from the perspective of whether extra parameters are involved: (1) extra-parameter methods, freeze all of the original parameters of an LLM and insert a set of learnable parameters to optimize the model input or model layers suach as adapter tuning [Houlsby et al., 2019] , prompt tuning [Lester et al., 2021] and prefix tuning [Li and Liang, 2021]; (2) intra-parameter methods freeze most of the original parameters of an LLM and only tune a small number of parameters of the LLM such as LoRA [Hu et al., 2022]. Furthermore, He et al. [2022] present a unified framework that establishes connections between PEFT methods and Zhu et al. [2024] formally identify and investigate asymmetry in the roles of low-rank adapter matrices in LoRA fine-tuning.